# Accelerating Transformer Pre-Training with 2:4 Sparsity

### v1

@@ 5.1 Fast Computation of Transposable Masks
We aim to find such a mask matrix $\boldsymbol{\mathbf{M}}\in\{0,1\}^{r\times q}$ for every $\boldsymbol{\mathbf{W}}\in\mathbb{R}^{r\times q}$ in the FFN layer that 1) each adjoining $4\times 4$ block contains 8 non-zero positions; each row and column in the block occupies 2 non-zero elements exactly (see in Section A.2); 2) $\max_{\boldsymbol{\mathbf{M}}}\|\boldsymbol{\mathbf{M}}\odot\boldsymbol{\mathbf{W}}\|_{1}$. Then $\boldsymbol{\mathbf{M}}$ would be our targeting transposable mask.
As described in Eq. (5), both a transposable mask itself and its transposition conform to the format of 2:4 sparsity. Previous 2-approximation algorithm (Hubara et al., 2021) consists of two steps: sort elements, and pick elements out of the array. They claim that the procedure has less computational complexity. However, in practice, the sorting and picking process contains too many jumps in its control flow, and may be fatal to modern GPU architecture. To make full use of the GPUs’ parallel computation capability (SIMD and SIMT), we convert the transposable mask-search process into a convolution operation which traverse all the masks to obtain the optimal one in three steps:
@@ 3. $index=\operatorname{argmax}(out,dim=2)$
5.2 Acceleration of Gated Activation Functions
  • 1)
  • 2)
  • 3)
Different from the dense model, where output activations are row-major matrices, in FST, the output activations are column-major matrices. This property results in the third step being extremely time-consuming if conventionally $\boldsymbol{\mathbf{Z}}$ is accessed along the row dimension. To illustrate, Figure 5 shows that in a column-major matrix $\boldsymbol{\mathbf{Z}}$, accessing along the column accords with memory layout. Thus, adjacent elements loaded into the GPU cache can be probably hit. By contrast, accessing along the row does not fully utilize the efficiency of GPU cache. In light of this, we carefully implement a GEGLU kernel where elements are accessed along the column dimension. In this way, GEGLU is performed 5 times faster than the naive counterpart (Table 3).
@@ 5.2 Acceleration of Gated Activation Functions
Figure 5: left: adapted method; right: intuitive method
@@ 5.3 Other Implementation Details
Reducing updating frequency. In fact, we find that a 2:4 mask doesn’t change a lot after one optimization step, and it is not necessary to update a mask frequently. For the sake of efficiency, we update the transposable masks of weights every $l$ optimizer steps. We usually take $l=40$ in practice.
Utilities. For 2:4-spMMs, we use CUTLASS (Thakkar et al., 2023) and cuSPARSElt (Mishra et al., 2021). Other GPU kernels are implemented in Triton, including a transposable mask search kernel, a pruning kernel, an MVUE kernel, a GEGLU kernel, and a masked decay kernel.
@@ 6 Experiments
method loss avg score cola mnli mnliextra mrpc qnli qqp rte sst2 stsb
@@
params method val loss avg score cola mnli mrpc qnli qqp rte sst2 stsb wnli
@@
Of note, we use $n$ to denote the length of sequences, $d$ to denote the input and output dimensions of each transformer block, $d_{ff}$ to denote the inner dimensions of the FFNs in each transformer block, $h$ to denote the number of heads, and $N$ to denote the micro-batch size on each device.
@@ 6.1 Accuracy Results
To investigate the effect of different 2:4 sparse training methods, we pre-train a sparse BERT-base model on the C4 dataset using two sparse training methods: STEP (Lu et al., 2023) and Bi-Mask (Zhang et al., 2023). Besides, we also pre-train a dense BERT-base and a ‘Half’ BERT-base for comparison. Of note, ‘Half’ denotes a smaller yet still dense BERT-base model. To create Half model, we simply reduce the $d_{ff}$ of each FFN layer in the original BERT-base by half while maintaining the original value of $d$. Theoretically, this adjustment halves the floating operations (FLOPs) of the original BERT-base as well. Except for the FFN layers, the shapes of the rest layers remain unaltered.
@@
Table 6: Experimental Results for Transformer-base and DeiT-tiny.
model matric Dense Ours
Transformer- base avg epoch loss 4.555 4.651
val loss 3.979 3.989
test BLEU 26.19 $26.38$
Deit-tiny test acc1 72.91 70.4
test acc5 91.57 $90.14$
test loss 1.19 1.303
Table 7: Experimental results of masked decay, MVUE, and dense fine-tuning (FT) with BERT-Base. For decay term, we use both techniques in Section 4.2 and 4.3.
decay mvue dense FT loss avg score
@@
Besides, we pre-train the GPT-2 series models with the proposed method. Table 5 shows that our method achieves lossless average scores compared with dense baselines for model sizes of 124M, 350M, and 775M. Similarly, Transformer-base and DeiT-tiny trained with our method also reach comparable results as dense training, as seen in Table 6.
Ablation study. We aim to investigate the effect of masked decay, MVUE and dense fine-tuning introduced in Section 4.2, 3.2, and 4.4. The 16-layer BERT-base is used for ablation study. Results in Table 7 show that: 1) The dense fine-tuning procedure helps to improve accuracy by 2 points at most on GLUE; 2) MVUE leads to insignificant, controllable accuracy loss; 3) By combining all these techniques together, 2:4 sparse training for transformers achieves comparable accuracy results as dense training.
6.2 Speedup Results
The training acceleration techniques proposed in Section 5 are evaluated using GPT-2 models and on RTX3090 GPUs. FP16 mixed precision training is used on all models. The practical speedups of a single FFN layer, a single transformer block, and the entire network, compared to their respective dense counterparts, are reported. All the measured datum contain both forward and backward propagation.
Feed-forward network layers. For a single FFN layer, we fix $n=2048$ and change $d$. Results in Figure 6 show that a FFN layer can be accelerated up to  1.5x faster than its corresponding dense layer. The speedup effect fades as the batch size becomes larger due to GPU memory bottleneck.
Transformer block. We measure the acceleration ratio of a transformer block when $n=512,1024,2048$. Results in Figure 6 show that in most cases, a transformer block can be accelerated to 1.2x faster via 2:4 sparsity.
@@ 6.2 Speedup Results
124M 16 1.1
350M 16 1.15
774M 5 1.2
End-to-end acceleration. Finally, we test the practical speedups of training GPT-2 models. Results in Table 8 show that our training method conducts up to 1.2x faster than the dense training on a single RTX3090.
@@ 7 Conclusions
In this study, we are the first to propose accelerating the pre-training of transformers by 2:4 sparsity. We analyze the limitations of previous 2:4 training methods, including the impropriety in choosing positions and determining values of the masked decay factor, speed bottleneck incurred by computing transposable masks and gated activation functions. We propose a series of techniques to tackle them. Our training method is validated on DeiT, BERT, Transformer and GPT-2 models. In particular, we have attained 1.2x end-to-end training acceleration for the GPT-2 774M model without losing its accuracy.
8 Broader Impact
@@
Examples of row-wise 2:4 sparse matrix and column-wise 2:4 sparse matrix are depicted in Figure 7.
Figure 7: Row-Wise and Column-Wise 2:4 Sparsity.
A.2 2:4 Transposable Masks
Example of a transposable mask with 2 blocks is depicted in Figure 8.
Figure 8: A Transposable Mask With 2 Blocks.
A.3 Memory Layout
The memory layout of different type of matrix multiplications are listed in Table 9, which explains why output activations and activation gradients are column-major matrices in FST.
Table 9: Memory layout of $\boldsymbol{\mathbf{M}}\boldsymbol{\mathbf{N}}$. Here $S$ denotes that the matrix is in row-wise 2:4 sparsity, $R$ denotes row-major dense matrix, and $C$ denotes column-major dense matrix.
@@ B.1 Forward and Backward Pass of a Linear Layer
Figure 9: 2:4 sparse training iteration for a layer on a single micro-batch.

### v3

@@ 5.1 Fast Computation of Transposable Masks
We aim to find such a mask matrix $\boldsymbol{\mathbf{M}}\in\{0,1\}^{r\times q}$ for every $\boldsymbol{\mathbf{W}}\in\mathbb{R}^{r\times q}$ in the FFN layer that 1) each adjoining $4\times 4$ block contains 8 non-zero positions; each row and column in the block occupies 2 non-zero elements exactly; 2) $\max_{\boldsymbol{\mathbf{M}}}\|\boldsymbol{\mathbf{M}}\odot\boldsymbol{\mathbf{W}}\|_{1}$. Then $\boldsymbol{\mathbf{M}}$ would be our targeting transposable mask.
As described in Equation 5, both a transposable mask itself and its transposition conform to the format of 2:4 sparsity. Previous 2-approximation algorithm (Hubara et al., 2021) consists of two steps: sort elements, and pick elements out of the array. They claim that the procedure has less computational complexity. However, in practice, the sorting and picking process contains too many jumps in its control flow, and may be fatal to modern GPU architecture. To make full use of the GPUs’ parallel computation capability (SIMD and SIMT), we convert the transposable mask-search process into a convolution operation which traverse all the masks to obtain the optimal one in three steps:
@@ 3. $index=\operatorname{argmax}(out,dim=2)$
Figure 6: left: adapted method; right: intuitive method
@@
5.2 Acceleration of Gated Activation Functions
  • 1)
  • 2)
  • 3)
Different from dense model, where output activations are row-major matrices, in FST, the output activations are column-major; see Section A.2. This property results in the third step being extremely time-consuming if conventionally $\boldsymbol{\mathbf{Z}}$ is accessed along the row dimension. To illustrate, Figure 6 shows that in a column-major matrix $\boldsymbol{\mathbf{Z}}$, accessing along the column accords with array layout. Thus, adjacent elements loaded into the GPU cache can be probably hit. By contrast, accessing along the row does not fully utilize the efficiency of GPU cache. In light of this, we carefully implement a GEGLU kernel where elements are accessed along the column dimension. In this way, GEGLU is performed 5 times faster than the naive counterpart; see Table 4.
@@ 5.3 Other Implementation Details
Reducing Updating Frequency
We find that a 2:4 mask doesn’t change a lot after one optimization step, and it is not necessary to update a mask frequently. For the sake of efficiency, we update the transposable masks of weights every $l$ optimizer steps. We usually take $l=40$ in practice.
Utilities
For 2:4-spMMs, we use CUTLASS (Thakkar et al., 2023). Other GPU kernels are implemented in Triton, including transposable mask search kernel, pruning kernel, MVUE kernel, GEGLU kernel, and masked decay kernel.
@@ 6 Experiments
Method Loss Avg score CoLA MNLI mnliextra MRPC QNLI QQP RTE SST-2 STS-B
@@
Params Method Val loss Avg Score CoLA MNLI MRPC QNLI QQP RTE SST-2 STS-B WNLI
@@
Table 7: SQuAD scores on GPT-2 models.
Params Method EM F1
124M Dense 67.6 78.8
Ours $67.5$ $78.5$
350M Dense 73.2 83.6
Ours $71.9$ $82.4$
774M Dense 74.3 84.9
Ours $74.3$ $84.6$
Table 8: Experimental results for DeiT.
Size Method Acc@1 Acc@5
DeiT-tiny Original [^2] Results reported in the original paper; see https://github.com/facebookresearch/deit/blob/main/README_deit.md. 72.2 91.1
Dense [^3] DeiT-base dense model using the original recipe. 72.9 91.6
Ours $70.4$ $90.1$
DeiT-small Original 79.9 90.5
Dense 79.9 94.5
Bi-Mask 77.6 -
Ours 79.2 94.8
DeiT-base Original 81.8 95.6
Dense 81.0 95.0
Ours $81.3$ $95.4$
Table 9: Experimental results for Transformer-base.
Method Avg epoch loss Test BLEU Val BLEU Val loss
Dense 4.558 26.15 26.56 3.982
Half 4.659 26.12 26.36 4.041
STEP 4.692 25.27 25.85 4.082
Ours $4.649$ $26.48$ $26.78$ $3.977$
@@
Of note, we use $n$ to denote the length of sequences, $d$ to denote the input and output dimensions of each transformer block, $d_{ff}$ to denote the inner dimensions of the FFNs in each transformer block, $h$ to denote the number of heads, and $N$ to denote the micro-batch size on each device. The pre-training and evaluation scripts are publicly available at https://github.com/thu-ml/2by4-pretrain-acc-examples.
@@ 6.1 Accuracy Results
To investigate the effect of different 2:4 sparse training methods, we pre-train a sparse BERT-base model on the C4 dataset using two sparse training methods: STEP (Lu et al., 2023) and Bi-Mask (Zhang et al., 2023). Besides, we also pre-train a dense BERT-base and a ‘Half’ BERT-base for comparison. Of note, ‘Half’ denotes a smaller yet still dense BERT-base model. To create Half model, we simply reduce the $d_{ff}$ of each FFN layer in the original BERT-base by half while maintaining the original value of $d$. Theoretically, this adjustment halves the floating operations (FLOPs) of the original FFN layer as well. Except for the FFN layers, the shapes of the rest layers remain unaltered.
@@
Table 10: Experimental results of masked decay, MVUE, and dense fine-tuning (FT) with BERT-Base. For decay term, we use both techniques in Sections 4.2 and 4.3.
Masked decay MVUE Dense FT Loss Avg score
@@
124M 16 1.18
350M 8 1.2
774M 4 1.21
Besides, we pre-train GPT-2 models with proposed methods. Table 6 and 7 shows that our method for model sizes of 124M, 350M, 775M and 1558M achieves lossless scores compared with dense baselines. Similarly, DeiT and Transformer-base trained with our method also reach comparable results to dense training; see Table 8 and 9. For GPT-2 and BERT, the training loss curves are sketched in Appendix C.
Ablation Study
We aim to investigate the effect of masked decay, MVUE and dense fine-tuning introduced in Section 4.2, 3.2, and 4.4. The 16-layer BERT-base is used for ablation study. Results in Table 10 show that: 1) The dense fine-tuning procedure helps to improve accuracy on GLUE by 2 points at most ; 2) MVUE leads to insignificant, controllable accuracy loss; 3) By combining all these techniques together, 2:4 sparse training for transformers achieves comparable accuracy results as dense training.
6.2 Speedup Results
The training acceleration techniques proposed in Section 5 are evaluated using GPT-2 models and RTX3090 GPUs. FP16 mixed precision training is used on all models. The practical speedups of a single FFN layer, a single transformer block, and the entire network, compared to their respective dense counterparts, are reported. All the measured datum contain both forward and backward propagation.
Feed-forward Network Layers
For a single FFN layer, we fix $n=2048$ and change $d$. Results in Figure 7 show that a FFN layer can be accelerated up to 1.7x faster than its corresponding dense layer.
Transformer Block
We measure the acceleration ratio of a transformer block when $n=512,1024,2048$. Results in Figure 7 show that in most cases, a transformer block can be accelerated to 1.3x faster via 2:4 sparsity. To illustrate this, a detailed profile result is given in Appendix D.
End-to-end Acceleration
Finally, we test the practical speedups of training GPT-2 models. Results in Table 11 show that our training method conducts up to 1.2x faster than the dense training on a single RTX3090.
@@ 7 Conclusions
In this study, we are the first to propose accelerating the pre-training of transformers by 2:4 sparsity. We analyze the limitations of previous 2:4 training methods, including the impropriety in choosing positions and determining values of the masked decay factor, speed bottleneck incurred by computing transposable masks and gated activation functions. We propose a series of techniques to tackle them. Our training method is validated on DeiT, BERT, Transformer-base and GPT-2 models. In particular, we have attained 1.2x end-to-end training acceleration for the GPT-2 774M model without losing its accuracy.
Acknowledgements
We would like to thank Ziteng Wang, Bingrui Li and Haocheng Xi for valuable discussions and help on the training large transformers. This work was supported by the National Key Research and Development Program of China (No. 2021ZD0110502), NSFC Projects (Nos. 62376131, 62061136001, 62106123, 62076147, U19A2081, 61972224), Tsinghua Institute for Guo Qiang, and the High Performance Computing Center, Tsinghua University. J.Z is also supported by the XPlorer Prize.
Impact Statement
@@ Acknowledgements
Examples of row-wise, column-wise and transposable 2:4 sparse matrix are shown in Figure 8. Note that transposable 2:4 sparsity aligns with both row-wise and column-wise 2:4 sparsity.
Figure 8: Row-wise 2:4, column-wise and transposable 2:4 sparse matrix.
A.2 Array Layout
The array layout of different types of matrix multiplications are listed in Table 12, which explains why output activations and activation gradients are column-major matrices in FST.
Table 12: Array layout of $\boldsymbol{\mathbf{M}}\boldsymbol{\mathbf{N}}$. Here $S$ denotes that the matrix is in row-wise 2:4 sparsity, $R$ denotes row-major dense matrix, and $C$ denotes column-major dense matrix.
@@ Appendix B Workflow
Figure 9: 2:4 sparse training iteration for a layer on a single batch.
@@ Appendix C Training Loss Curve
Appendix D Profiling result
To explain how we reach 1.3x block speedup, we profile our code and break down the time costs as shown in the table below; see Table 13.
Table 13: Time costs of each part of our network and the dense model in one iteration per layer. $m$ denotes the accumulation steps over micro batches. Our method is evaluated on GPT-2, with batch size 16, sequence length 1024, embedding dimension 1024 and heads number 16.
Dense (ms/exec) Sparse (ms/exec) Acceleration ratio $S$ Frequency (exec/iter)
FFN Linear FWD GEMM 12173.8 7305.78 1.666324472 -
BWD GEMM 23295 14080.82 1.654378083 -
MVUE+prune 0 171.4 - -
Total 23295 14252.22 1.634482207 -
Total 35468.8 21558 1.645273216 -
Others [^4] All functions in FFN except linear layers, i.e., activation function and dropout. FWD 167 118.17 - -
BWD 65.5 20.03 - -
Total 232.5 138.2 - -
Total FWD 12340.8 7423.95 1.662295678 -
BWD 23360.5 14272.25 1.636777663 -
Total 35701.3 21696.2 1.645509352 -
Others FWD 6874.3 7090.55 - -
BWD 13920.7 14117.45 - -
Total 20795 21208 - -
Total FWD 19215.1 14514.5 1.323855455 -
BWD 37281.2 28389.7 1.313194574 -
Total 56496.3 42904.2 1.316801152 -
Masked decay 0 45.2 - $\frac{1}{m}$
Prune weights 0 320.3 - $\frac{1}{m}$
Transposable mask search 0 634.8 - $\frac{1}{40m}$